Find the limit superior and the limit inferior of the following sequences






a) {(1 +






1






𝑛






)






𝑛+1






}






b) {






(βˆ’1)






𝑛






𝑛2






}

1
Expert's answer
2021-12-20T11:28:54-0500

Solution:

Assume the sequences are as follows:

(a) "\{a_n\}=\{\frac{1}{n}+(-1)^n\}""|a_n|=|\frac{1}{n}+(-1)^n|\leq2" for all integers "n\geq1". Hence "\{a_n\}" is bounded.

The first few terms are "\{0,\frac{3}{2},\frac{-2}{3},\frac{5}{4},\frac{-4}{5},\frac{7}{6},\cdots\}". The subsequence "\{a_{2n}\}=\{\frac{3}{2},\frac{5}{4},\frac{7}{6},\cdots\}" converges to 1 and the subsequence "\{a_{2n-1}\}=\{0,\frac{-2}{3},\frac{-4}{3},\cdots\}" converges to -1. Hence "S=\{-1,1\}"

"\liminf a_n=\inf(S)=-1\\
\limsup a_n=\sup(S)=1"

(b)

"x_{n}=(-1)^{n} \frac{1+n}{n}"

Since "\frac{1+n}{n}=1+\frac{1}{n} \geq 0" for any "n \in \mathbb{N}" we have

"\lim _{n \rightarrow \infty} \sup x_{n}=\lim _{k \rightarrow \infty} x_{2 k}=1"

"\limsup _{n \rightarrow \infty} x_{n}=\lim _{k \rightarrow \infty} x_{2 k}=1"

and

"\liminf _{n \rightarrow \infty} x_{n}=\lim _{k \rightarrow \infty} x_{2 k+1}=-1"